# Boson-fermion pairing and condensation in two-dimensional Bose-Fermi mixtures

### v1

@@ I Introduction
Mixed systems of bosons and fermions are at the base of our understanding of the physical world. They feature in the Standard Model of elementary-particle physics as gauge and matter fields, respectively, in high density quark matter as diquarks interacting with unpaired quarks [1], in nuclear matter in the interacting boson-fermion model for atomic nuclei with an odd number of protons or neutrons [2], and as boson-mediated interactions in soft [3] as well as condensed [4] matter. Ultra-cold atomic gases provide an exceptionally versatile platform for the simulation of quantum matter [5]. In the case of Bose-Fermi (BF) atomic mixtures, several realizations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] have permitted to explore a wide range of phenomena like phase separation [22], polarons [13, 17, 23, 26, 31], dual superfluidity [15, 16, 19, 20], collective excitations [34, 35], mediated interactions [28, 36, 37, 38], and Feshbach molecules [6, 8, 10, 25, 30, 33]
In a parallel fashion, BF mixtures have also been realized in optical semiconductor micro-cavities [40, 41, 42] thanks to the recent advent of atomically thin transition metal dichalcogenides [43, 44]. These materials intrinsically offer a remarkable opportunity to investigate the physics in two dimensions (2D) where topological phases play a crucial role [45, 46].
@@
However, the unique flexibility of atomic gases is unrivaled when the need arises to manipulate dimensionality and interaction strength at the same time. Confinement induced resonances accomplish this purpose by allowing for the realization of a dimensional crossover, starting from 3D confinement down to (quasi-)2D one. If the energy level spacing of the 3D trapping system (or equivalently the confinement length) is kept much larger than the energy (shorter than the length) scales of the intervening scattering processes, the system becomes effectively 2D and the ensuing 2D scattering length can be manipulated by varying either the scattering length associated to the original 3D Feshbach resonance and/or the confinement length, yet ensuring the quasi-2D condition above remains satisfied [47, 48, 49]. This situation is particularly apt for BF mixtures, for which one would need to adjust independently the BF and the boson-boson (BB) interactions owing to the mechanical stability of the assembly being conditioned primarily by that of the Bose component [50]. Having at one’s disposal two independent degrees of freedom in the 3D scattering length and in the confinement length, one can then correspondingly tune the effective BF and BB interactions (even though in a highly non-linear fashion [47]) and simultaneously analyze the stability of different phases of the mixture across the whole resonant regime.
@@
Under such circumstances, a recent work has shown that two-dimensional weakly bound fermionic molecules formed in a BF mixture with attractive BF and repulsive BB interactions are expected to be collisionally stable and should exhibit a strong p-wave mutual attraction [51]. In fact three-body recombination, which has historically plagued these mixed systems in three dimensions, is concomitantly avoided because Efimovian processes are suppressed in 2D.
@@
$P$-wave superfluidity is intensely pursued in both atomic based [52] and semiconductor based [53, 45] mixtures, as it opens the way to the exploration of topologically protected states, including Majorana modes and non-Abelian vortex excitations [54, 55], hence offering a route for the implementation of fault-tolerant quantum computation [56, 57, 58, 59, 60, 61].
@@
In this work, we study a 2D BF mixture in homogeneous space with boson and fermion densities spanning from the bosonic impurity limit (${n_{\mathrm{B}}}\ll{n_{\mathrm{F}}}$) to the matched density regime (${n_{\mathrm{B}}}={n_{\mathrm{F}}}$). Similarly to the analyses carried out in 3D [73, 77, 81, 80, 83, 88], we are here interested in the competition between boson condensation and BF pairing in mixtures with an attractive and tunable BF interaction at zero temperature. We focus on the case with concentration of bosons $x={n_{\mathrm{B}}}/{n_{\mathrm{F}}}\leq 1$, such that a full competition between pairing and condensation is allowed. The BB interaction is considered to be repulsive for mechanical stability reasons.
@@
Our main results are as follows. (i) We found that the condensate is progressively depleted as the BF attraction increases, in analogy with the 3D case. However, while in 3D the condensate vanishes beyond a critical coupling strength, in 2D the condensate does not exactly vanish but rather becomes exponentially small with the coupling strength. (ii) Similarly to the 3D case, we uncover a nearly universal behavior of the condensate fraction, the bosonic momentum distribution, and Tan’s contact parameter with respect to the boson concentration. (iii) BF pairs form for sufficiently strong BF attraction. Quite generally, these composite fermions are hybridized states between a pure molecular bound state and an unpaired state made of a fermionic atom and a boson belonging to the condensate. We found that these hybridized states produce quasi-particles with rather peculiar Fermi liquid features. (iv) The boson momentum distribution function show quite a rich behavior when varying the BF coupling strength and boson concentration, which could in principle be tested in future experiments with BF mixtures.
@@
The article is organized as follows. In Sec. II we introduce the model Hamiltonian, illustrate the diagrammatic theory of choice and lay out the full set of thermodynamic equations to solve. In Sec. III we explore the strong-coupling limit for the BF coupling parameter. In this limit, a set of semi-analytical and rather simple equations can be derived, allowing us to obtain crucial insights on the microscopic physics and phenomenology of a strongly attractive BF mixture in 2D. In Sec. IV we present our numerical results for the boson and fermion momentum distributions, corresponding chemical potentials, condensate density and Tan’s contact parameter. In Sec. V we discuss virtues and limitations of our approach and outline perspectives for future work.
II Formalism
II.1 The system Hamiltonian
@@
We consider a mixture of bosons (B) and single-component fermions (F) with interspecies attraction in a two-dimensional homogeneous setting, and focus on the zero temperature limit in which bosons may condense even in 2D [89].
@@
where $\psi^{\dagger}_{s}$ and $\psi_{s}$ respectively create and destroy a particle of mass $m_{s}$ and chemical potential $\mu_{s}$, with $s=\mathrm{B,F}$ indicating the bosonic or fermionic nature. We consider two assemblies with densities ${n_{\mathrm{B}}}\leq{n_{\mathrm{F}}}$, relative concentration $x={n_{\mathrm{B}}}/{n_{\mathrm{F}}}$ and equal masses ${m_{\mathrm{B}}}={m_{\mathrm{F}}}$ and adopt the convention $\hbar=k_{\mathrm{B}}=1$ throughout this paper.
@@
The BF attractive contact interaction needs to be regularized by expressing the bare strength $v_{0}^{\rm BF}$ in terms of the scattering length $a_{\mathrm{BF}}$. This is achieved by the equation
[…]
relating the bare coupling strength $v_{0}^{\mathrm{BF}}$ with the boson-fermion binding energy $\varepsilon_{0}=1/(2m_{r}a_{\mathrm{BF}}^{2})$ of the associated two-body problem in vacuum, where $a_{\mathrm{BF}}$ is the (2D) BF scattering length and $m_{r}={m_{\mathrm{B}}}{m_{\mathrm{F}}}/({m_{\mathrm{B}}}+{m_{\mathrm{F}}})$ is the reduced mass of the boson-fermion system. The ultraviolet-divergent integral on the right-hand side of Eq. (2) compensates analogous divergencies occurring in many-body diagrammatic perturbation theory.
Based on physical intuition, and in analogy with the 3D case, one expects the following picture when the BF coupling strength is varied. In the weak coupling limit, bosons are expected to be almost fully condensed (compatibly with a quantum depletion determined by $\eta_{\mathrm{B}}$) and fermions to fill up a Fermi sphere with chemical potential renormalized at the mean-field level. In the strong coupling limit, bosons are instead expected to be expelled out of the condensate and form molecular fermions that fill up a Fermi sphere. We will discuss below to what extent this simple picture continues to hold also in two dimensions.
Figure 1: Feynman’s diagrams for the $T$-matrix $\Gamma(P)$ in the normal phase (a) and $T(P)$ in the condensed phase (b). Full lines correspond to bare boson (B) and fermion (F) Green’s functions, dashed lines to bare BF interactions and zigzag lines to condensate factors $\sqrt{n_{0}}$.
II.2 Many-body $T$-matrix
The diagrammatic many-body approach adopted in the present work relies on previous works on BF mixtures in 3D, both above [73] and below [88] the BEC transition temperature, based on the non-self-consistent $T$-matrix (ladder) approximation. In these works, the $T$-matrix approach was shown to recover the well-known polaron-molecule transition of an impurity in a polarized Fermi gas as the coupling strength is increased and to predict a universal behavior of the condensate fraction as a function of the coupling strength with respect to the boson concentration. Such outcomes were recently confirmed in a break-through experiment on 3D BF mixtures, where for the first time the regime of double degenerate BF mixtures with matching densities was reached [33]. There the authors observed the existence of a quantum critical point at the polaron-to-molecule phase transition and confirmed the universal character of the condensate fraction with respect to the boson concentration, predicted in [88] some time before. Based on these solid and sound results, we keep the same diagrammatic choice of [88] and extend it to the 2D case.
@@ II.2 Many-body $T$-matrix
with $G_{\mathrm{B}}^{0}$ and $G_{\mathrm{F}}^{0}$ the bare Green’s functions for bosons and fermions
@@
where $\xi_{\mathbf{k}}^{s}=k^{2}/2m_{s}-\mu_{s}$.
Following Appendix A, one obtains
@@
where $T_{2}$ is the off-shell two-body $T$-matrix in vacuum and the contribution $I_{F}({\mathbf{P}},\Omega)$ stems from the presence of the fermionic medium (see Appendix A for details). Note that the present diagrammatic approach corresponds to the zero-temperature limit of the Matsubara formalism at finite temperature. We will thus work on the imaginary frequency axis rather than on the real frequency one (as instead is standard practice at $T=0$ [92]). This choice is particularly useful to avoid singularities of the Green’s function on the real frequency axis.
We now consider the condensed phase, for which additional diagrams are to be taken into account [88]. These are shown in Fig. 1(b) with zigzag lines representing condensate factors: here additional fluctuation processes are introduced by allowing the BF pairs to break up and exchange bosons with the condensate. The resulting Bethe-Salpeter equation for the $T$-matrix in the condensed phase $T({\mathbf{P}},\Omega)$ thus reads
@@
Viewing $T({\mathbf{P}},\Omega)$ and $\Gamma({\mathbf{P}},\Omega)$ as dressed and undressed pair propagators respectively, the above equation can be interpreted as a Dyson’s equation for the pair propagator $T({\mathbf{P}},\Omega)$, with the related self-energy insertion $n_{0}G_{\mathrm{F}}^{0}({\mathbf{P}},\Omega)$ representing the exchange of bosons with the condensate.
II.3 Boson and fermion self-energies
@@
The bosonic self-energy is made of two terms: the interaction with fermions and that with other bosons. The former is shown in Fig. 2(a), yielding the expression
@@
Note here that we adopt a sign convention such that the $T$-matrix $T({\mathbf{P}},\Omega)$ reduces to the $T$-matrix in vacuum [93].
@@
$\begin{split}&\Sigma_{\mathrm{B}}\left(\textbf{k},\omega\right)=\Sigma_{11}+\\ &\int\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\int\frac{d\Omega}{2\pi}T\left(\textbf{P},\Omega\right)G_{\mathrm{F}}^{0}\left(\textbf{P}-\textbf{k},\Omega-\omega\right)e^{i\Omega 0^{+}}.\end{split}$ (10)
The diagrams contributing to the fermionic self-energy are of two types, since they originate from closing the $T$-matrix $T({\mathbf{P}},\Omega)$ either with a normal boson propagator or with two condensate insertions, as Fig. 2(b) shows. However, since the latter choice gives rise to an improper self-energy contribution, to avoid double counting the $T$-matrix in the normal phase $\Gamma({\mathbf{P}},\Omega)$ is used instead, thus yielding
$\begin{split}&\Sigma_{\mathrm{F}}\left(\textbf{k},\omega\right)=n_{0}\Gamma\left(\textbf{k},\omega\right)-\\ &\int\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\int\frac{d\Omega}{2\pi}T\left(\textbf{P},\Omega\right)G_{\mathrm{B}}^{0}\left(\textbf{P}-\textbf{k},\Omega-\omega\right)e^{i\Omega 0^{+}}.\end{split}$ (11)
II.4 Green’s functions, momentum distributions, and densities

### v2

@@ 1 Introduction
Mixed systems of bosons and fermions are at the base of our understanding of the physical world. They feature in the Standard Model of elementary-particle physics as gauge and matter fields, respectively, in high-density quark matter as diquarks interacting with unpaired quarks [1], in nuclear matter in the interacting boson-fermion model for atomic nuclei with an odd number of protons or neutrons [2], and as boson-mediated interactions in soft [3] as well as condensed [4] matter. Ultra-cold atomic gases provide an exceptionally versatile platform for the simulation of quantum matter [5]. In the case of Bose-Fermi (BF) atomic mixtures, several realizations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] have permitted to explore a wide range of phenomena like phase separation [22], polarons [13, 17, 23, 26, 31], dual superfluidity [15, 16, 19, 20], collective excitations [34, 39], mediated interactions [28, 35, 36, 37], and Feshbach molecules [6, 8, 10, 25, 30, 33].
In a parallel fashion, BF mixtures have also been realized in semiconductor nanostructures thanks to the recent advent of atomically thin transition metal dichalcogenides (TMD) [40, 41, 42, 43, 44, 45]. These materials intrinsically offer a remarkable opportunity to investigate the physics in two dimensions (2D) where topological phases play a crucial role [46, 47]. The so-called spin-valley locking, originating from the strong spin-orbit coupling of the transition metal atoms, allows for a complete distinguishability between electrons (and/or holes) populating the two spin-polarized valleys at Brillouin points $K$ and $K^{\prime}$ [48]. Exciton formation and charge doping are therefore two distinct processes affecting separate assemblies of electrons (and/or holes). In this scenario, the production of a BF mixture in a two-dimensional solid-state setting becomes therefore possible, with bosons corresponding to tightly bound excitons and fermions to doped charges. The recent detection of electrically tunable two-dimensional Feshbach resonances in TMD bilayers has put this target on more solid ground [43, 49].
@@
However, the unique flexibility of atomic gases is unrivaled when the need arises to manipulate dimensionality and interaction strength at the same time. Confinement-induced resonances accomplish this purpose by allowing for the realization of a dimensional crossover, starting from 3D confinement down to (quasi-)2D one. If the energy level spacing of the 3D trapping system (or equivalently the confinement length) is kept much larger than the energy (shorter than the length) scales of the intervening scattering processes, the system becomes effectively 2D and the ensuing 2D scattering length can be manipulated by varying either the scattering length associated with the original 3D Feshbach resonance and/or the confinement length, yet ensuring the quasi-2D condition above remains satisfied [50, 51, 52]. This situation is particularly apt for BF mixtures, for which one would need to adjust independently the BF and the boson-boson (BB) interactions owing to the mechanical stability of the assembly being conditioned primarily by that of the Bose component [53]. Having at one’s disposal two independent degrees of freedom in the 3D scattering length and the confinement length, one can then correspondingly tune the effective BF and BB interactions (even though in a highly non-linear fashion [50]) and simultaneously analyze the stability of different phases of the mixture across the whole resonant regime.
@@
Under such circumstances, a recent work has shown that two-dimensional weakly bound fermionic molecules formed in a BF mixture with attractive BF and repulsive BB interactions are expected to be collisionally stable and should exhibit a strong $p$-wave mutual attraction [54]. In fact, three-body recombination, which has historically plagued these mixed systems in three dimensions, is concomitantly avoided because Efimovian processes are suppressed in 2D.
@@
$P$-wave superfluidity is intensely pursued in both atomic-based [55] and semiconductor-based [56, 46] mixtures, as it opens the way to the exploration of topologically protected states, including Majorana modes and non-Abelian vortex excitations [57, 58], hence offering a route for the implementation of fault-tolerant quantum computation [59, 60, 61, 62, 63, 64].
@@
In this work, we study a 2D BF mixture in homogeneous space with boson and fermion densities spanning from the bosonic impurity limit (${n_{\mathrm{B}}}\ll{n_{\mathrm{F}}}$) to the matched density regime (${n_{\mathrm{B}}}={n_{\mathrm{F}}}$). Similarly to the analyses carried out in 3D [76, 80, 84, 83, 86, 91], we are here interested in the competition between boson condensation and BF pairing in mixtures with an attractive and tunable BF interaction at zero temperature. For this reason, we consider concentrations of bosons $x={n_{\mathrm{B}}}/{n_{\mathrm{F}}}\leq 1$, such that a full competition between pairing and condensation is allowed. The BB interaction (when considered) is taken repulsive for mechanical stability reasons. Numerical calculations focus on the case of equal boson and fermion masses in order to reduce the computational effort. However, to facilitate possible future studies, the theoretical formalism is developed for generic boson and fermion masses $m_{\rm B}$ and $m_{\rm F}$.
@@
Our main results are as follows. (i) We find that the condensate is progressively depleted as the BF attraction increases, in analogy with the 3D case. However, while in 3D the condensate vanishes beyond a critical coupling strength, in 2D the condensate does not exactly vanish but rather becomes exponentially small with the coupling strength. (ii) Similarly to the 3D case, we uncover a nearly universal behavior of the condensate fraction, the bosonic momentum distribution, and Tan’s contact parameter with respect to the boson concentration. (iii) BF pairs form for sufficiently strong BF attraction. Quite generally, these composite fermions are hybridized states between a pure molecular bound state and an unpaired state made of a fermionic atom and a boson belonging to the condensate. We find that these hybridized states produce quasi-particles with rather peculiar Fermi liquid features. (iv) The boson momentum distribution function shows quite a rich behavior when varying the BF coupling strength and boson concentration, which could in principle be tested in future experiments with BF mixtures.
@@
The article is organized as follows. In Sec. 2 we introduce the model Hamiltonian, illustrate the diagrammatic theory of choice and lay out the full set of thermodynamic equations to solve. In Sec. 3 we explore the strong-coupling limit for the BF coupling parameter. In this limit, a set of semi-analytical and rather simple equations can be derived, allowing us to obtain crucial insights into the microscopic physics and phenomenology of a strongly attractive BF mixture in 2D. In Sec. 4 we present our numerical results for the boson and fermion momentum distributions, corresponding chemical potentials, condensate density and Tan’s contact parameter. In Sec. 5 we discuss the virtues and limitations of our approach and outline perspectives for future work.
2 Formalism
2.1 The system Hamiltonian
@@
We consider a mixture of bosons (B) and single-component fermions (F) with interspecies attraction in a two-dimensional homogeneous setting and focus on the zero temperature limit in which bosons may condense even in 2D [92].
@@
where $\psi^{\dagger}_{s}$ and $\psi_{s}$ respectively create and destroy a particle of mass $m_{s}$ and chemical potential $\mu_{s}$, with $s=\mathrm{B,F}$ indicating the bosonic or fermionic nature. We set $\hbar=1$ throughout this paper.
@@
[…]
On physical grounds, one expects the following picture when the BF coupling strength is varied. In the weak-coupling limit, bosons are expected to be almost fully condensed, with a small depletion of the condensate density $n_{0}$ determined just by the weak BB repulsion [94] for $g\to-\infty$. Fermions instead fill up a Fermi sphere. In the strong-coupling limit, bosons are instead expected to be expelled out of the condensate and form molecular fermions that fill up a Fermi sphere. We will discuss below to what extent, within the present approach, this simple picture continues to hold also in two dimensions.
Figure 1: Feynman’s diagrams for (a) the particle-particle ladder $\Gamma(P)$ and (b) the $T$-matrix in the condensed phase $T(P)$. Full lines correspond to bare boson (B) and fermion (F) Green’s functions, dashed lines to bare BF interactions and zigzag lines to condensate factors $\sqrt{n_{0}}$.
2.2 Many-body $T$-matrix
The diagrammatic many-body approach adopted in the present work relies on previous works on BF mixtures in 3D, both above [76] and below [91] the BEC transition temperature, based on the non-self-consistent $T$-matrix (ladder) approximation. In these works, such approximation was shown to recover the well-known polaron-to-molecule transition of an impurity in a polarized Fermi gas as the coupling strength is increased, and to predict a universal behavior of the condensate fraction with respect to the boson concentration (from $x\to 0$ to $x=1$) as the coupling strength is varied. Such outcomes were recently confirmed in a breakthrough experiment on 3D BF mixtures, where for the first time the regime of double degenerate BF mixtures with matching densities was reached [33]. There, the authors observed the existence of a quantum phase transition between a polaronic condensed phase and a molecular phase and, even for nearly matched densities, confirmed the predictions of Ref. [91] on the universal character of the condensate fraction with respect to the boson concentration. These successes in 3D, motivate us to keep the same diagrammatic choice of [91] and extend it to the 2D case.
A warning is however in order. In the limiting case of a single impurity in a Fermi gas, it has been shown that the present approach misses the polaron-to-molecule transition in 2D [97], which is instead found by more sophisticated approaches [98, 99]. We should thus bear in mind that such a shortcoming of the present approximation, which is due to an overestimate of the atom-molecule repulsion in the strong-coupling limit of the theory, may extend also at finite boson concentration $x$, in particular in the limit $x\to 0$ that approaches the single impurity case.
Let us see now the rationale behind the choice of diagrams made in [91], which we wish to extend to 2D. Since the pioneering work by Nozières and Schmitt-Rink [100] on the related problem of the BCS-BEC crossover in two-component Fermi gases, it is well-known that ladder diagrams are able to capture pairing (molecular) correlations in the normal phase (see also [101] for a review). It is only after the inclusion of this class of diagrams that the superfluid critical temperature recovers the Bose-Einstein condensation temperature in the strong-coupling limit of the BCS-BEC crossover. In addition, for the same contact potential we are considering, ladder diagrams provide the leading self-energy in the weak-coupling limit of the BCS-BEC crossover [102]. They thus provide a sensible scheme to describe the whole BCS-BEC crossover, even in the intermediate coupling region in which fully controlled approximations are not available.
[…]
where $G_{\mathrm{B}}^{0}$ and $G_{\mathrm{F}}^{0}$ are bare Green’s functions for bosons and fermions
@@ 2.2 Many-body $T$-matrix
with $\xi_{\mathbf{k}}^{s}=k^{2}/2m_{s}-\mu_{s}$.
Integrating over $\omega$ and solving for $\Gamma\left(\textbf{P},\Omega\right)$ in Eq. (3), one obtains
[…]
whereby the $\Theta$ functions appearing in the integrand correspond to the Fermi and Bose distributions in the zero temperature limit. In particular, for the concentrations $x\leq 1$ and the range of BF and BB coupling strengths considered in this work, ${\mu_{\mathrm{B}}}$ is always negative, hence $\Theta(-\xi_{\mathbf{k}}^{\mathrm{B}})=0$.
One then obtains (see Appendix A for details)
@@
where $T_{2}$ is the off-shell two-body $T$-matrix in vacuum and the contribution $I_{\rm F}({\mathbf{P}},\Omega)$ stems from the presence of the fermionic medium.
Note that the present diagrammatic approach corresponds to the zero-temperature limit of the Matsubara formalism at finite temperature. We will thus work on the imaginary frequency axis rather than on the real frequency one (as instead it is standard practice at $T=0$ [103]). This choice is particularly useful to avoid singularities of the Green’s function on the real frequency axis.
We now consider the condensed phase. In this case, one has to take into account the possibility that fermions repeatedly interact also with condensed bosons, besides non-condensed bosons. By summing all possible combinations of the repeated scattering of fermions with condensed or non-condensed bosons one obtains the many-body $T$-matrix in the condensed phase described by the Feynman diagram of Fig. 1(b). The resulting Bethe-Salpeter equation for the many-body $T$-matrix in the condensed phase $T({\mathbf{P}},\Omega)$ thus reads
@@
2.3 Boson and fermion self-energies
@@
The bosonic self-energy is made of two terms: originating, respectively, from BF and BB interactions. For the former, we proceed like in the corresponding problem in the BCS-BEC crossover: the self-energy of one species is obtained by closing the $T$-matrix by a propagator of the other species.
This is shown in Fig. 2(a), yielding the expression
@@
where the factor $e^{i\Omega 0^{+}}$ resulting from Feynman’s rules [103] ensures convergence of the frequency integral. Note here that we adopt a sign convention such that the many-body $T$-matrix $T({\mathbf{P}},\Omega)$ reduces to the two-body $T$-matrix in vacuum [104].
@@
$\Sigma_{\mathrm{B}}\left(\textbf{k},\omega\right)=\Sigma_{11}+\Sigma_{\mathrm{BF}}\left(\textbf{k},\omega\right).$ (11)
For the fermionic self-energy in the condensed phase, the bosonic line closing the many-body $T$-matrix $T({\bf P},\Omega)$ can be either a condensed line or a non-condensed one. In the first case, however, the many-body $T$-matrix in the condensed phase $T({\bf P},\Omega)$ needs to be replaced by $\Gamma({\bf P},\Omega)$ in order for the self-energy to be irreducible (and thus avoiding a double-counting of diagrams when inserting the self-energy in the Dyson equation). The fermionic self-energy is thus described by the Feynman diagrams of Fig. 2(b), corresponding to the expression
$\Sigma_{\mathrm{F}}\left(\textbf{k},\omega\right)=n_{0}\Gamma\left(\textbf{k},\omega\right)-\int\!\!\!\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\int\!\!\!\frac{d\Omega}{2\pi}T\left(\textbf{P},\Omega\right)G_{\mathrm{B}}^{0}\left(\textbf{P}-\textbf{k},\Omega-\omega\right)e^{i\Omega 0^{+}}.$ (12)
2.4 Green’s functions, momentum distributions, and densities